# Communication to Completion: Modeling Collaborative Workflows with Intelligent Multi-Agent Communication

### v1

@@ Abstract
Teamwork in workspace for complex tasks requires diverse communication strategies, but current multi-agent LLM systems lack systematic frameworks for task oriented communication. We introduce Communication to Completion (C2C), a scalable framework that addresses this gap through two key innovations: (1) the Alignment Factor (AF), a novel metric quantifying agent task alignment that directly impacts work efficiency, and (2) a Sequential Action Framework that integrates stepwise execution with intelligent communication decisions. C2C enables agents to make cost aware communication choices, dynamically improving task understanding through targeted interactions. We evaluated C2C on realistic coding workflows across three complexity tiers and team sizes from 5 to 17 agents, comparing against no communication and fixed steps baselines. The results show that C2C reduces the task completion time by about 40% with acceptable communication costs. The framework completes all tasks successfully in standard configurations and maintains effectiveness at scale. C2C establishes both a theoretical foundation for measuring communication effectiveness in multi-agent systems and a practical framework for complex collaborative tasks.
@@ 1 Introduction
Figure 1: Communication to Completion architecture: a five-layer stack simulation engine
Modern organizations run on teams: people plan, divide, and verify work under deadlines while juggling synchronous meetings, asynchronous chats/emails, and limited attention. Extending this paradigm to AI, coordinating teams of LLM agents to execute complex, long-horizon tasks (e.g., software development) remains a grand challenge (Park et al., 2023; Hong et al., 2024; Qian et al., 2023). Effective collaboration hinges on communication: too much creates coordination overhead; too little yields misalignment and rework.
Recent multi-agent frameworks demonstrate that structured dialogues and role specialization can transform general purpose LLMs into cooperative problem solvers across software engineering, data analysis, and decision support (Wu et al., 2024; Hong et al., 2024; Qian et al., 2023; Li et al., 2023). Beyond mere tool use, conversation itself decomposes tasks, aligns partial knowledge, and arbitrates proposals, often outperforming single agent pipelines on complex problems (de Souza Santos and Ralph, 2022; Khan et al., 2024; Yao et al., 2023a; Besta et al., 2024). Yet many systems still schedule or trigger communication via fixed or purely reactive heuristics that ignore the evolving trade-off between coordination cost and task progress (Wu et al., 2024; Chen et al., 2023).
Meanwhile, empirical studies of software teams warn that coordination costs (especially meetings), can dominate delivery time in modern distributed work, with a post-2020 shift toward heavier synchronous load and changing patterns across chat, email, and meetings (Mok et al., 2023; Yang et al., 2022). These observations motivate treating communication as a first class resource to be scheduled, routed, and evaluated rather than as a byproduct of agent reasoning, thus leading to a central open question: how can agents learn to communicate strategically to maximize collaborative benefit while minimizing coordination costs?
In this work, we address the above question by casting multi-agent collaboration as cost communication and optimizing it accordingly. Our contributions are:
  • We introduce Communication to Completion (C2C). A scalable framework models when, whom, and how to communicate as part of execution and instantiate it with a Sequential Action Framework (SAF) that synchronizes agent actions into discrete timesteps for deterministic, reproducible collaboration.
  • We define Alignment Factor (AF), a metric that quantifies each agent’s task understanding and directly modulates work efficiency. Communications update AF, and AF in turn scales work rate, linking conversation to progress and yielding interpretable traces.
  • We demonstrate through comprehensive experiments spanning complexities and team sizes, C2C lowers completion time and improves work efficiency compared to baselines.
@@ 2 Related Work
Multi-Agent Collaboration Frameworks
LLM agents are increasingly organized as teams rather than monoliths. Social simulation work demonstrates emergent, human-like collective behavior (e.g., the “Smallville” agents of Park et al. (2023) and the recent large scale AgentSociety simulation (Piao et al., 2025)). For goal oriented collaboration—particularly software development—frameworks such as MetaGPT (Hong et al., 2024) and ChatDev (Qian et al., 2023) impose predefined roles and SOPs that mirror corporate workflows. AutoGen (Wu et al., 2024) offers a flexible substrate for composing conversable agents, while AgentVerse (Chen et al., 2023) and OpenAgents (Xie et al., 2023) provide extensible environments for building and deploying general purpose agents. Outside pure research prototypes, platforms like OpenHands operationalize multi/single agent development loops in realistic toolchains (Wang et al., 2025). In parallel, role-playing and deliberation structures further organize collaboration within conversations (Li et al., 2023; Yao et al., 2023a; Besta et al., 2024), and new evaluations emphasize interactive, multi-environment testing and repository-level search and planning (Liu et al., 2023a; Antoniades et al., 2024). Our work differs in elevating communication from a procedural step to an optimizable resource, explicitly linking when/what to communicate with collaborative progress via an alignment driven mechanism.
Communication in Multi-Agent Systems
Communication has long been central in multi-agent learning. Classic MARL studies learned protocols end-to-end under partial observability (Foerster et al., 2016; Sukhbaatar et al., 2016). Recent advances focus on efficient or sparse communication to reduce overhead, e.g., decentralized scheduling and model-based message estimation (Han et al., 2023; Liu et al., 2023c; Kim et al., 2019; Singh et al., 2018). In LLM agent settings, natural-language messaging carries high-level, semantic content, and methods increasingly structure agent talk through reasoning–acting loops (ReAct) (Yao et al., 2023b), self improvement with textual feedback (Reflexion) (Shinn et al., 2023), and iterative self-refinement (Madaan et al., 2023). Multi-agent debate further uses conversational arbitration to improve correctness and planning (Du et al., 2023). Yet many systems still trigger communication on fixed or purely reactive heuristics. We instead propose an alignment driven policy that decides when, whom and how to communicate based on a measurable, task grounded alignment signal.
Quantifying Coordination and Alignment
Coordination costs are a foundational concern in collaborative systems (Malone, 1988; Brooks Jr, 1995). However, evaluations of LLM multi agents often emphasize final outcomes (e.g., pass rates on benchmarks) over process efficiency. In software engineering, SWE-bench exposes the gap between realistic issue resolution and model capabilities (Jimenez et al., 2023), while SWE-agent (Yang et al., 2024) and OpenHands (Wang et al., 2025) reveal the importance of tool use and orchestration—but still provide limited real-time signals about team alignment. Complementary efforts argue for richer process traces and interactive environments to analyze coordination choices (Liu et al., 2023a; Antoniades et al., 2024), and work on LLM-as-judge warns that automated evaluators may introduce bias at the process level (Liu et al., 2023b; Chen et al., 2024; Tan et al., 2024). We contribute the Alignment Factor (AF) as a dynamic, real-time proxy for shared task understanding that both modulates work efficiency and directly informs agents’ communication decisions, thereby connecting conversational actions to tangible task outcomes.
@@ 3 C2C Framework
Communication to Completion is a time-stepped framework for deciding when, whom, and how to communicate in multi-agent collaboration. Figure 2 provides an overview of its mechanism using a software task. We organize the discussion as follows: §3.1 introduces the Sequential Action Framework that ensures deterministic execution, §3.2 describes the Alignment Factor mechanism for quantifying task understanding, §3.3 details how agents make communication decisions, §3.4 explains hierarchical task decomposition and tracking, and §3.5 presents the intention-based agent decision making that integrates these components.
3.1 Sequential Action Framework
Concurrent multi-agent execution often introduces temporal ambiguity. The Sequential Action Framework (SAF) constrains each agent to exactly one action per timestep, yielding deterministic state transitions and reproducible analysis.
@@ 3.1 Sequential Action Framework
SAF defines four actions capturing common collaborative behavior that capture common collaborative behavior: work (task execution with time cost), communicate (compose/send a message), reply (respond to an incoming message), and meeting (synchronous group discussion). Each action is temporally bounded and commits a single transition.
@@
Forward-only Delivery.
@@ 3.2 Alignment Factor
The Alignment Factor (AF) measures an agent’s task-specific understanding and directly modulates effective productivity. Unlike static capability scores, AF evolves through communication. For agent $i$ on task $j$, $AF_{i,j}\in[0.01,1.00]$ denotes comprehension quality. Values near $0.01$ indicate limited understanding; values near $1.00$ indicate mastery. We initialize $AF_{i,j}{=}0.3$ to reflect the initial understanding.
When agent $i$ receives reply about task $j$ from agent $k$, AF is updated based on the communication’s impact on task understanding. An LLM evaluates the message content, agent $i$’s current task context, and the relevance of the received information to determine the new alignment:
@@
where $\Delta_{\text{eval}}\in[0,0.5]$ is computed by the LLM based on: (i) how well the message addresses the agent’s knowledge gaps, (ii) the relevance of the information to the specific task requirements, and (iii) the clarity and actionability of the guidance provided. Different message types naturally yield different impacts, help requests addressing critical blockers typically produce larger alignment gains.
@@
A low AF impedes an agent’s progress, creating a natural incentive for it to seek clarification and coordination before investing substantial effort.
@@ 3.3 Agent Communication Decisions
Agents autonomously decide when, with whom, and how to communicate based on their current task state and situational awareness, without relying on fixed thresholds or rigid scheduling. For clarity, we decompose this decision process into four key components: initiating the communication, selecting the recipients, choosing the appropriate channel, and composing the message content.
@@
Agents may initiate communication in response to various task situations: encountering technical difficulties (help_request), facing unclear requirements (need_clarification), identifying coordination needs with team members (meeting_invite), or reaching milestones that warrant status updates (progress_update). The decision to communicate emerges from the agent’s assessment of its current alignment and task complexity.
@@
When initiating communication, agents consider multiple factors in selecting recipients: skill relevance to the issue, existing task dependencies, and historical interaction. The selection process reflects realistic collaboration patterns observed in human teams.
Channel Selection
Agents choose among three communication channels. Chat suits quick questions and simple clarifications with rapid response times. Email handles detailed explanations with longer response windows. Meetings address complex coordination needs requiring synchronous discussion among multiple participants.
@@
Communication content is generated based on the agent’s current context, including task description, encountered difficulties, current alignment factor, and relevant technical details. Messages are composed to be informative and actionable, providing recipients with sufficient context to offer effective assistance.
@@ 3.5 Intention-Based Agent Decision Making
Given this context, the agent uses an LLM to carefully reason about the current situation and generate an intention, a high-level decision about what action to take next. The generated intention is then translated into a concrete action within the SAF framework:
  • Work: Continue task execution with effectiveness modulated by current AF.
  • Help: Compose the message identifying difficulties and required expertise (HELP_REQUEST).
  • Clarification: Request additional task details or requirement explanations (NEED_CLARIFICATION).
  • Coordination: Propose synchronous discussion for complex collaboration needs (MEETING_INVITE).
  • Reporting: Share subtasks progress with manager (PROGRESS_REPORT).
@@ 4.1 Experimental Setup
Metric Complexity No Communication Fixed Steps C2C
Task Completion Rate (%) Simple 100 100 100
Medium 100 100 100
Complex 100 100 100
Avg Completion Time (hours) Simple 7 5 5.5
Medium 20 14.75 13
Complex 33.5 36.25 24.75
Communication Cost Simple 2.03 1.94
Medium 2.75 3.26
Complex 8.12 7.02
Alignment Score (AF) Simple 0.3 0.55 0.51
Medium 0.3 0.59 0.68
Complex 0.3 0.53 0.55
Efficiency Simple 1.14 1.6 1.45
Medium 1.20 1.63 1.85
Complex 1.19 1.10 1.62
Table 1: Comparison of three policies for a 1M+4W team across task complexities (Simple=8h, Medium=24h, Complex=40h). We report Task Completion Rate (%), Avg Completion Time (hours), Communication Cost (total communication time), Alignment Score (mean AF over the run; higher indicates better shared understanding), and Efficiency == completed work / total time. C2C achieves the best time and efficiency on Medium and Complex with comparable communication cost. Bold marks the best value within each row; "-" denotes not applicable.
To evaluate the performance of C2C on realistic tasks, we implement experiments on software engineering workflows across three complexity tiers: Simple (8 hours, basic SWE operations), Medium (24 hours, API integration with authentication), and Complex (40 hours, full stack user authentication system); see Appendix A for details. Each task requires diverse skills including front-end development, back-end systems, database management, security implementation, and testing. Tasks are decomposed into several subtasks according to task complexity and team size with explicit skill requirements and dependencies, mirroring real-world software development scenarios.
Team Configurations
We systematically vary team sizes from 5 to 17 agents: 1M+4W (5 agents), 1M+8W (9 agents), and 1M+16W (17 agents), where M denotes manager and W denotes worker. Additionally, we evaluate multi-task scenarios with 1M+8W handling 2 concurrent tasks to assess workload distribution and task interference effects. Each agent is assigned 2–4 complementary skills from a skill pool, including skills like frontend, backend, and database.
Models and Implementation
All agents are powered by GPT-4o with temperature 0.7 for human like decision making. The Sequential Action Framework operates with 0.25 hour time steps over a maximum of 160 simulation steps (40 hours). Communication costs are calculated using realistic time models: email drafting (9 minutes base + content length), chat messages (3 minutes base), meetings (30 minutes minimum + preparation time + number of participants). We implement thread-depth limits (maximum 3 reply rounds) to prevent infinite communication loops.
Baselines
We compare C2C against two systematic baselines: (1) No Communication: agents work independently with task assignments but no communication, representing traditional parallel processing approaches; (2) Fixed Steps: agents communicate every 16 steps, simulating conventional project management practices. Both baselines use identical task decomposition and skill matching but lack C2C’s intelligent communication strategies.
Metrics
Following our evaluation framework, we report: task completion rate (percentage of tasks completed within time budget), average completion time (hours to finish successful tasks), communication cost (total time spent on communication activities), alignment score (average agent task alignment factor), and efficiency (ratio of productive work to total time investment). Each configuration is evaluated across all three task complexity tiers.
4.2 Main Results
As shown in Table˜1, which details the results for a 1M+4W team, C2C consistently demonstrates superior performance in key efficiency metrics, although all methods achieved a 100% task completion rate. The primary advantage of C2C is most evident in task completion time. On Complex tasks, C2C finishes in 24.75 hours, significantly faster than both No Communication (33.5 hours) and Fixed Steps (36.25 hours). This trend holds for Medium tasks, where C2C is also the fastest.
@@ 4.2 Main Results
This time saving translates to higher efficiency. For Complex tasks, C2C’s efficiency score is 1.62, substantially outperforming Fixed Steps (1.10) and No Communication (1.19). Similarly, C2C achieves the highest alignment score on Medium (0.68) and Complex (0.55) tasks, confirming that its communication strategy effectively enhances agent understanding. While communication costs are comparable across strategies, C2C utilizes its communication budget more effectively to achieve better outcomes.
Figure 3 illustrates the dynamics of the alignment factor mechanism. Alignment scores start low as agents have minimal task understanding, then improve through strategic communications. Help requests and meetings yield the strongest alignment improvements, while progress updates are modest. The communication heatmap reveals the manager centric communication pattern.

### v2

@@ Abstract
Multi-agent LLM systems have demonstrated impressive capabilities in complex collaborative tasks, yet most frameworks treat communication as instantaneous and free, overlooking a fundamental constraint in real world teamwork, collaboration cost. We propose a scalable framework implemented via Communication to Completion (C2C), which explicitly models communication as a constrained resource with realistic temporal costs. We introduce the Alignment Factor (AF), a dynamic metric inspired by Shared Mental Models, to quantify the link between task understanding and work efficiency. Through experiments on 15 software engineering workflows spanning three complexity tiers and team sizes from 5 to 17 agents, we demonstrate that cost-aware strategies achieve over 40% higher efficiency compared to unconstrained interaction. Our analysis reveals emergent coordination patterns: agents naturally adopt manager centric hub-and-spoke topologies, strategically escalate from asynchronous to synchronous channels based on complexity, and prioritize high value help requests. These patterns remain consistent across multiple frontier models (GPT-5.2, Claude Sonnet 4.5, Gemini 2.5 Pro). This study moves beyond simple agent construction, offering a theoretical foundation for quantifying and optimizing the dynamics of collaboration in future digital workplaces.
@@ 1 Introduction
Figure 1: Overview of the C2C simulation framework. Agents autonomously select between execution and communication channels by weighing the Alignment Factor (AF) against modeled latencies, forming a closed-loop system where communication is treated as a constrained resource.
Effective teamwork in software development hinges on strategic communication. Teams must continuously balance the value of information sharing against its costs: synchronous meetings consume collective time, asynchronous messages introduce delays, and frequent interruptions fragment attention (Mok et al., 2023; Yang et al., 2022). Understanding when, with whom and how to communicate remains a central challenge in organizational research (Marlow et al., 2018).
However, systematically studying these dynamics in real teams is notoriously difficult due to the inability to control variables or replay scenarios. Recent advances in large language models (LLMs) offer a new approach to simulate workplace behaviors. Just as Park et al. (2023) demonstrated that LLM agents can exhibit believable human-like social behavior in simulated communities, we propose using multi-agent LLM systems as a controlled simulation environment for investigating workplace collaboration patterns.
Despite this potential, current multi-agent frameworks are not well-suited for realistic workplace simulation. Systems such as AutoGen and MetaGPT typically abstract away communication costs, effectively assuming instantaneous information exchange (Wu et al., 2024; Hong et al., 2024). As a result, agents can exchange unlimited tokens without explicitly modeling temporal latency or cognitive costs (e.g., context switching) inherent in real workflows. Consequently, these frameworks cannot capture the fundamental trade-off between alignment and coordination cost, and may encourage excessive communication that would be unsustainable for human teams.
To bridge this gap, we introduce Communication to Completion (C2C), a framework designed to simulate collaborative workflows under realistic constraints. C2C incorporates a discrete event simulation engine that explicitly models communication costs, enabling communication to be treated as a scarce resource to be optimized. Within this environment, we propose the Alignment Factor (AF), a dynamic metric inspired by the concept of Shared Mental Models (Cannon-Bowers et al., 1993), to quantify how effectively agents build and maintain shared task understanding over time. We use AF both as an evaluation metric and as a control signal to guide when additional communication is worth its cost.
Our contributions are as follows:
  • We establish a computational sandbox for multi-agent collaboration that moves beyond instantaneous communication assumptions, enabling the study of agent behaviors under realistic temporal and cognitive constraints.
  • We introduce the Alignment Factor as a quantifiable metric for shared task understanding. We show that optimizing for AF enables agents to dynamically balance the benefits of information sharing against the costs of coordination.
  • Through experiments on complex software engineering tasks, we demonstrate that cost aware communication strategies outperform unconstrained interaction in efficiency. We further analyze emergent coordination patterns, showing how constraints induce more structured communication topologies.
@@ 2 Related Work
2.1 Multi-Agent Collaboration Frameworks
Recent research on LLM multi-agent systems broadly spans two lines of work: task oriented execution frameworks and behavioral simulations. Task oriented frameworks such as MetaGPT (Hong et al., 2024), ChatDev (Qian et al., 2023), and AutoGen (Wu et al., 2024) employ role structures, Standard Operating Procedures (SOPs), or conversable agents to solve complex engineering tasks. While effective for end-to-end task completion, these systems typically do not explicitly model communication as a constrained resource: interaction is commonly treated as low friction message passing, without representing temporal latency or cognitive overhead that shape real world teamwork. Conversely, behavioral simulations such as Generative Agents (Park et al., 2023) and AgentSociety (Piao et al., 2025) focus on reproducing believable human social dynamics and daily routines, but are less tied to goal directed engineering workflows. Our work bridges these directions by adopting the simulation perspective while targeting task-oriented collaboration: we explicitly model coordination constraints and costs to enable controlled study of communication strategies under resource limitations.
2.2 Communication Costs and Constraints
Communication overhead is a foundational concern in organizational theory and collaborative systems, and has also been widely studied in multi-agent reinforcement learning (MARL). In MARL, agents learn to selectively communicate or gate messages to reduce bandwidth usage and coordination overhead (Singh et al., 2018; Kim et al., 2019). In LLM-based multi-agent systems, communication is predominantly natural language and is often treated as part of the reasoning process rather than an explicit resource. Although recent work structures agent interaction through protocols such as debate and other deliberative exchanges (Du et al., 2023; Khan et al., 2024), mainstream frameworks still rarely represent costs such as asynchronous delay or context switch overhead. Empirical studies in software engineering further suggest that coordination costs (e.g., synchronous meetings and scheduling friction) can dominate delivery time in distributed teams (Mok et al., 2023; Yang et al., 2022). Our framework explicitly models such costs via a discrete event simulation engine, distinguishing our approach from systems that view communication primarily as a byproduct of reasoning.
2.3 Shared Mental Models and Alignment
Effective teamwork relies on Shared Mental Models (SMMs), the overlapping knowledge structures held by team members that support coordinated action (Cannon-Bowers et al., 1993; Mathieu et al., 2000). In human teams, misalignment can lead to “process loss”, where effort is diverted to resolving misunderstandings rather than executing tasks. For LLM agents, assessing alignment remains challenging. Most evaluations emphasize outcome correctness (e.g., pass@1 on SWE-bench (Jimenez et al., 2023; Yang et al., 2024)) rather than process quality. While recent benchmarks and analyses examine interaction trajectories (Liu et al., 2023a; Yang et al., 2024), they do not provide a dynamic measure of consensus during collaboration. We introduce the Alignment Factor (AF) as a computational proxy for SMM, enabling quantitative measurement of how communication updates shared context and how evolving alignment relates to downstream execution efficiency.
@@ 3 C2C Framework
Communication to Completion is a simulation framework for deciding when, with whom, and how to communicate in multi-agent collaboration. Figure 2 provides an overview of the mechanism. We organize the discussion as follows: §3.1 introduces the discrete event engine that ensures deterministic execution, §3.2 describes the Alignment Factor for quantifying task understanding, §3.3 details how agents make cost-aware communication decisions, and §3.4 explains hierarchical task decomposition.
3.1 Discrete Event Simulation Engine
Concurrent multi-agent execution often introduces temporal ambiguity. To ensure reproducibility, we employ a discrete event simulation engine that constrains each agent to exactly one action per timestep, yielding deterministic state transitions.
@@ 3.1 Discrete Event Simulation Engine
The engine defines four actions capturing common collaborative behavior: work (task execution with time cost), communicate (compose/send a message), reply (respond to an incoming message), and meeting (synchronous group discussion). Each action is temporally bounded and commits a single transition.
@@
Forward-only Delivery
@@ 3.2 Alignment Factor
The Alignment Factor (AF) serves as a proxy for the Shared Mental Model (Cannon-Bowers et al., 1993), measuring an agent’s task specific understanding. Unlike static capability scores, AF evolves through communication. For agent $i$ on task $j$, $AF_{i,j}\in[0.01,1.00]$ denotes comprehension quality. We initialize $AF_{i,j}{=}0.3$ to reflect the initial ambiguity.
When agent $i$ receives information, the AF is updated based on the communication’s quality. To rigorously measure this without agent bias, the update is computed by the simulation environment (using ground-truth context) rather than the agent itself:
@@
where $\Delta_{\text{eval}}\in[0,0.5]$ is computed based on: (i) knowledge gap resolution, (ii) relevance to task requirements, and (iii) actionability.
@@
A low AF impedes progress, creating a natural incentive to seek clarification despite the communication cost.
@@ 3.3 Agent Communication Decisions
Agents autonomously decide communication strategies based on situational awareness. Crucially, this decision process is cost aware, balancing the need for alignment against simulated channel constraints.
@@
Agents may initiate communication in response to task situations such as technical difficulties (help_request) or uncoordinated dependencies (meeting_invite). The decision emerges from the agent’s assessment of its current AF versus the time cost of interaction.
@@
Agents select recipients based on skill relevance and historical interaction, reflecting realistic collaboration patterns.
Channel Selection and Costs
Agents operate on a utility maximization principle. We model the utility $U_{i,ch}$ for agent $i$ choosing channel $ch$ as:
$U_{i,ch}=w_{1}\cdot\mathcal{S}(D_{task},B_{ch})-w_{2}\cdot\mathcal{C}(T_{lat},N)+\rho_{i,ch}$ (4)
where:
  • Task Suitability ($\mathcal{S}$): Measures how well the channel’s bandwidth $B_{ch}$ matches the current task complexity $D_{task}$. Complex tasks yield higher utility in high bandwidth channels (Meetings), whereas simple tasks (e.g., status check) favor low ones.
  • Time Cost ($\mathcal{C}$): Penalizes transmission latency $T_{lat}$ and synchronization overhead for $N$ participants.
  • Agent Preference ($\rho_{i,ch}$): Captures agent specific biases, $\rho>0$ for preferred channels.
This composite formulation ensures that agents do not merely seek the fastest route, but dynamically select the most appropriate channel given the task’s nature and their role constraints.
@@
Communication content is generated based on the agent’s current context, including task description and alignment gaps. Messages are composed to be informative.
@@ 3.5 Intention-Based Agent Decision Making
Metric Complexity No Comm Fixed Steps Free Comm C2C (Ours)
Task Completion Rate (%) Simple $100$ $100$ $96\pm 8$ $100$
Medium $100$ $100$ $100$ $100$
Complex $100$ $100$ $100$ $100$
Avg Completion Time (hours) $\downarrow$ Simple $7.00$ $\mathbf{5.00\pm 0.3}$ $6.00\pm 0.4$ $5.50\pm 0.25$
Medium $20.0$ $14.75\pm 0.75$ $14.5\pm 1.1$ $\mathbf{13.00\pm 0.9}$
Complex $33.5$ $36.25\pm 3.25$ $28.0\pm 2.8$ $\mathbf{24.75\pm 1.5}$
Communication Cost (hours) $\downarrow$ Simple $2.03\pm 0.1$ $3.20\pm 0.6$ $\mathbf{1.94\pm 0.2}$
Medium $2.75\pm 0.2$ $5.40\pm 1.5$ $\mathbf{3.26\pm 0.2}$
Complex $8.12\pm 0.9$ $14.55\pm 3.2$ $\mathbf{7.02\pm 0.8}$
Alignment Score (AF) $\uparrow$ Simple $0.30$ $0.55\pm 0.05$ $\mathbf{0.60\pm 0.04}$ $0.51\pm 0.03$
Medium $0.30$ $0.59\pm 0.06$ $0.63\pm 0.05$ $\mathbf{0.68\pm 0.04}$
Complex $0.30$ $0.53\pm 0.08$ $\mathbf{0.65\pm 0.07}$ $0.55\pm 0.05$
Efficiency (Work/Time) $\uparrow$ Simple $1.14$ $\mathbf{1.60\pm 0.1}$ $1.33\pm 0.7$ $1.45\pm 0.07$
Medium $1.20$ $1.63\pm 0.1$ $1.66\pm 0.12$ $\mathbf{1.85\pm 0.13}$
Complex $1.19$ $1.10\pm 0.09$ $1.43\pm 0.14$ $\mathbf{1.62\pm 0.1}$
Table 1: Main results comparing collaboration policies on GPT-4o agents. Free Comm represents an unconstrained baseline where agents communicate without simulated delays, resulting in faster raw completion but high communication costs. Values are reported as mean $\pm$ standard deviation over 5 runs. C2C achieves the highest Efficiency (task progress normalized by total time investment) on Medium and Complex tasks, demonstrating the value of cost-aware strategy. ($\downarrow$: lower is better, $\uparrow$: higher is better).
@@
Given this context, the agent uses an LLM to generate an intention. The generated intention translates into actions within the simulation:
  • Work: Continue task execution (efficiency modulated by AF).
  • Help/Clarification: Request assistance (incurs communication cost).
  • Coordination: Propose meetings (high synchronization cost).
  • Reporting: Share progress.
@@ 4.1 Experimental Setup
To evaluate the performance of C2C on realistic tasks, we constructed a benchmark dataset consisting of 15 software engineering workflows across three complexity tiers (5 tasks per tier) to ensure robustness across different business domains (see Appendix A for full task list). For clarity in our detailed ablation studies, we report metrics from the representative task of each tier:
  • Simple (8 hours): Basic SWE operations (e.g., bug fixes, refactoring).
  • Medium (24 hours): Module integration (e.g., API integration with authentication).
  • Complex (40 hours): Full-stack system design (e.g., user authentication service).
Each task requires diverse skills including front-end development, back-end systems, database management, and testing. Tasks are decomposed into subtasks with explicit dependencies, mirroring real-world software development scenarios.
@@
Team Configurations
We systematically vary team sizes from 5 to 17 agents: 1M+4W (5 agents), 1M+8W (9 agents), and 1M+16W (17 agents), where M denotes manager and W denotes worker. Additionally, we evaluate multi-task scenarios with 1M+8W handling 2 concurrent tasks.
Models and Implementation
Agents are powered by GPT-4o (temperature 0.7). We also evaluate Claude 3.5 Sonnet to test generalization. The discrete event simulation engine operates with 0.25-hour time steps over a maximum of 160 steps. Communication costs model real world constraints: Email (9 min delay + drafting), Chat (3 min), and Meetings (30 min blocking cost).
Baselines
We compare C2C against three baselines: (1) No Communication: Agents work independently. (2) Fixed Steps: Agents communicate on a rigid schedule (every 16 steps). (3) Free Communication: An unconstrained baseline where agents communicate without simulated latency, representing idealized interaction.
Metrics
We report: task completion rate (percentage of tasks completed within time budget), average completion time (hours to finish successful tasks), communication cost (total time spent on communication activities), alignment score (average agent task alignment factor), and efficiency (ratio of productive work to total time investment). Each configuration is evaluated across all three task complexity tiers.
4.2 Main Results
Table 1 presents the performance comparison. While all methods achieve high completion rates, C2C consistently demonstrates superior Efficiency on Medium (1.85) and Complex (1.62) tasks. Comparison with the Free Communication baseline reveals the impact of communication overhead. Although Free Comm allows agents to communicate instantly, this leads to excessive interaction (e.g., 14.55 hours cost on Complex tasks vs. 7.02 hours for C2C). This chatter distracts agents from execution, effectively slowing down the overall progress. C2C, by weighing the cost of communication against alignment gains, finds an optimal balance, engaging in just enough coordination to maintain high Alignment Factor (AF) without incurring high costs.
To understand the dynamics behind this efficiency, Figure 3 analyzes the evolution of the Alignment Factor. As shown in Figure 3(a), alignment scores start low and improve progressively. Figure 3(b) highlights that high cost interactions like Meetings yield the largest alignment gains justifying their use for critical synchronization, whereas low cost messages provide smaller increments. The heatmap in Figure 3(c) confirms a manager coordination pattern, where the manager acts as a hub to distribute context efficiently.
4.3 Generalization across Models
To verify that our findings are not specific to GPT-4o, we evaluated C2C using three frontier models: GPT-5.2, Claude Sonnet 4.5, and Gemini 2.5 Pro. As shown in Table 2, the performance advantage of C2C remains consistent across all backbones. The Claude Sonnet 4.5 model, known for its specialized coding capabilities, achieves the highest raw performance. However, without constraints, it still incurs significant communication overhead in the Free Communication baseline (Efficiency 1.96). Applying the C2C framework boosts its efficiency to 2.40 (a 22% relative improvement). Similarly, GPT-5.2 sees an efficiency gain from 2.08 to 2.22, and Gemini 2.5 Pro from 1.88 to 2.09. These results confirm that while stronger models can execute tasks faster, they are not immune to coordination friction. C2C provides a universal optimization layer that scales with model capability, ensuring that high performance agents focus on execution rather than redundant communication.
Backbone Model Policy Time (h) $\downarrow$ Comm Cost (h) $\downarrow$ Efficiency $\uparrow$
GPT-5.2 Free Comm $11.50$ $4.20$ $2.08$
C2C (Ours) $\mathbf{10.80}$ $\mathbf{2.15}$ $\mathbf{2.22}$
Claude Sonnet 4.5 Free Comm $12.25$ $4.80$ $1.96$
C2C (Ours) $\mathbf{10.00}$ $\mathbf{2.40}$ $\mathbf{2.40}$
Gemini 2.5 Pro Free Comm $12.75$ $5.10$ $1.88$
C2C (Ours) $\mathbf{11.50}$ $\mathbf{2.65}$ $\mathbf{2.09}$
Table 2: Generalization analysis on Medium complexity tasks. Even with highly capable models, the unconstrained Free Comm baseline suffers from communication overhead. C2C consistently improves Efficiency by minimizing coordination costs, demonstrating that the framework’s value scales with model capability.
4.4 Scalability and Multi-Task Analysis
[…]